An iron rod 4m long & 0.5cm^2 in cross sectional area streches 1mm when a mass of 225kg is hong on it end. Extimate the stress and the strain experience by the iron rod?


1
Expert's answer
2021-05-24T15:58:07-0400

Gives

Length(l )=4m

"area(A)=0.5 cm^2=0.5\times10^{-4}m^2"

∆l=1mm="10^{-3}m"

Mass(M)=225kg

We know that

Stress "=\frac{F}{A}"

F=mg

"g=9.8m\/sec^2"

"F=225\times9.8N=2205N"



"Stress=\frac{2205}{0.5\times10^{-4}}=4.41\times10^7N/m^2"

Strain"=\frac{∆l}{l}"

Put value

Strain="\frac{1\times10^{-3}}{4}" ="2.5\times10^{-4}"